# Seq1F1B: Efficient Sequence-Level Pipeline Parallelism for Large Language Model Training

### v3

@@ 3.2 Seq1F1B
During the warm-up phase, the number of warm-up micro-batches of each worker $i$ is calculated according to Eq. 4, in which $k$ represents the number of splits in the sequence. This equation ensures that the last stage can perform a backward pass on the last sequence segment of the first batch when entering the steady phase, and the worker responsible for each stage performs one more forward pass than the worker responsible for the subsequent stage. Here, we construct a partially ordered queue $Q_{s}$, where each pop returns the tail sequence from the earliest batch that has enqueued. This satisfies the first-in-first-out principle in the batch dimension and the first-in-last-out principle in the sequence dimension. In each iteration of the warm-up phase, workers execute one forward pass and enqueue the corresponding hidden states onto $Q_{s}$.
In the steady phase, after each worker completes a forward pass, it dequeues from $Q_{s}$ and performs a backward pass on the dequeued hidden states, following the standard 1F1B process, except that the units for forward and backward passes become a sequence segment.
In the cooling-down phase, workers dequeue the remaining warm-up hidden states from $Q_{s}$ and perform backward passes sequentially.
@@
From the timeline shown in Figure 1, it is evident that the Seq1F1B schedule offers shorter execution time and significantly fewer bubbles compared to the original 1F1B schedule. Meanwhile, it can be clearly seen that each worker now has less memory consumption since the micro-sequence is smaller than the micro-batch. Another observation is that optimizations similar to the zero-bubble pipeline can also be applied to Seq1F1B by delaying the gradient computation associated with weights in the backward pass. We will discuss this in the following section.
3.3 Seq1F1B-I
Figure 3: Execution timeline for the zero-bubble-pipeline’s ZBH1 and Seq1F1B schedule intergrated with zero-bubble-pipeline’ ZBH1. Where each micro-batch is labeled with an ID and different colors distinguish the forward/backward/weight computation of different stages. OPT stands for optimizer step.
1F1B-I Narayanan et al. [2021a] achieves better efficiency by modifying the 1F1B schedule to support interleaved stages among workers. In 1F1B-I, each worker is assigned multiple stages. Suppose we have $P$ workers and $V$ stages $\{s_{1},s_{2},\ldots,s_{V}\}$ in our pipeline, where $V$ is a multiple of $P$. Each worker $i$ will handle $n$ stages $\{s_{i},s_{i+P},s_{i+2P},\ldots,s_{i+(n-1)P}\}$, where $n=\frac{V}{P}$. The number of warm-up micro-batches of each worker $i$ in 1F1B-I is given in Eq. 5.
$w_{i}=(P-i-1)\times 2+(n-1)\times P$ (5)
After completing $P$ iterations of forward/backward passes, each worker switches its context to the next stage it is responsible for. From the Figure 2, the above part shows a 1F1B-I pipeline with $P$ as 4 and $V$ as 8, in which each worker handles 2 stages. 1F1B-I’s schedule reduces the bubble ratio by interleaving stages among workers. However, this interleaving slightly increases memory consumption, as the number of warm-up micro-batches $w_{i}$ is greater compared to 1F1B.
Similar to 1F1B-I, Seq1F1B-I further modifies 1F1B-I to achieve sequence-level scheduling, as shown in bottom part of Figure 2. From the Figure 2, Seq1F1B-I effectively reduces pipeline bubbles and maintains less memory footprint of hidden state compared with 1F1B-I. Seq1F1B-I defines the number of warm-up micro-batches as in Eq. 6.
$w_{i}=(P-i-1)\times 2+(n-1)\times P+k-1$ (6)
in which $k$ represents the number of splits in the sequence. By using the partially ordered queue, Seq1F1B-I maintains strict order of forward/backward pass and ensures the consistent semantics of gradient updates. From the perspective of pipeline bubbles, Seq1F1B-I outperforms both Seq1F1B and 1F1B-I. In terms of memory demands, Seq1F1B-I requires slightly more memory than Seq1F1B but significantly less than 1F1B-I.
3.4 Integration with Zero-bubble-pipeline
From the illustration3, we can see Seq1F1B can integrate with ZB1P method and further reduce bubbles while reducing memory demands by splitting sequence. Such integration outperforms simple ZB1P in both memory demands and pipeline bubbles since sequence-level pipelines naturally have fewer bubbles. Furthermore, Seq1F1B can integrate with ZB2P and ZBV methods too. Theoretically, introducing a zero-bubble-pipeline to Seq1F1B should be more efficient. Even though, such a fine-grained handcraft schedule may have performance degradation under some settings. We hope our work inspires future work to solve this problem.
3.5 Workload Balance
In this section, we detail the strategy of sequence partition and workload balance consideration. Previous works, such as Li et al. [2021], have discussed strategies for sequence partitioning. To achieve efficient pipeline scheduling, it is crucial that the processing times for each subsequence are approximately equal to avoid pipeline bubbles. Based on this premise, we design a computation-wise partition strategy by estimating the FLOPs of sequences and constructing a theoretical solution aiming to make the FLOPs of all subsequences as closely as possible.
For a input sequence $S=(x_{1},x_{2},\cdots,x_{n})$, we devide it into $k$ segments $S=[S_{1},\cdots,S_{k}]$. Each segment having a length of $n_{i}$, where $\sum_{i=1}^{k}n_{i}=n$. We expect the computational amount of each segment to be roughly the same, that is
$\text{FLOPs}(S_{1})=\text{FLOPs}(S_{2})=\cdots=\text{FLOPs}(S_{k})={\frac{\text{FLOPs}(S)}{k}}.$ (7)
Specifically, we use the method proposed in Hoffmann et al. [2022] to estimate the FLOPs for each subsequence, as formulated in Eq. 8,
$\text{FLOPs}(S_{i})=2~{}n_{i}~{}P+2~{}L~{}n_{i}\left(\sum_{j=0}^{i}n_{j}\right)d,\forall i=1\dots k;\ \ \text{FLOPs}(S)=2~{}n~{}P+2~{}L~{}n^{2}d\end{aligned},$ (8)
in which, $L$ is a number of layers, $d$ is dimension of the model, and $P$ is the total number of parameters in the model. We have $k$ variables in Eq. 8 and $k$ equations in Eq. 7. Therefore, we can set up the equation to get the optimial segmentation.
4 Experiments
4.1 Experimental Settings
Table 1: Settings used in experiments for training LLMs.
@@ 4.1 Experimental Settings
Figure 4: Peak Memory consumption of tranining a series of models under varying sequence lengths and fixed batch settings. “X” means experiments ran out of memory. We take the maximum memory consumption between all workers for better clarification.
In experiments, we measure our methods and 1F1B and 1F1B-I under variable sequence lengths, different numbers of micro-batches, different numbers of GPUs, different pipeline parallel sizes and tensor parallel sizes. Compared methods are as follows:
  • Seq1F1B: Seq1F1B with computation-wise sequence partition strategy.
  • Seq1F1B-I: Seq1F1B with interleaved stages and computation-wise sequence partition strategy.
  • 1F1B/1F1B-I: 1F1B and 1F1B with interleaved stages in Megatron implementation.
  • Seq1F1B w/o cwp: Seq1F1B without computation-wise sequence partition strategy.
  • Seq1F1B-I w/o cwp: Seq1F1B-I without computation-wise sequence partition strategy.
All assessments are based on GPT model and model configuration are listed in Table 1. All experiments focus on long-sequence training since a lot of work has metioned the importance. For hyperparameter configurations, we set the number of sequence splits to four and each worker managing two stages in interleaved settings. Our implementation is based on the open-source Megatron-LM project Narayanan et al. [2021a] and ensures reproducibility. We adopts Megatron-V3Korthikanti et al. [2023]’s tensor parallelism in all experiments since it is necessary for long sequence training.
4.2 Main Results
Table 2: 2.7B GPT training experiments with pipeline parallel size of 8 under 8xA100 setting.
@@ 4.2 Main Results
Table 3: 7B GPT training experiments with pipeline parallel size of 4 and tensor parallel size of 8 under 32xA100 setting.
@@
Table 4: 13B GPT training experiments with pipeline parallel size of 4 and tensor parallel size of 8 under 32xA100 setting.
@@
Table 5: 30B GPT training experiments with pipeline parallel size of 8 and tensor parallel size of 8 under 64xA100 setting.
@@
In Figure  4, we compared the memory consumption of our method with that of 1F1B and 1F1B-I. As can be seen, our method consistently requires less memory across all settings, and notably, it can support training a 30B model on a 64xA100 cluster, which is impossible for the traditional combination of pipeline and tensor parallelism. Additionally, we recorded TFLOPS(teraFLOPS) per GPU in our experiments to measure the hardware utilization of different methods. From the Table 2, 3, 4 and 5, our method Seq1F1B outperforms 1F1B and 1F1B-I under almost all settings in both training throughput and teraFLOPS.
However, as observed in Table 3,4,5, the Seq1F1B-I may have a performance degradation under multi-node settings. This could be due to the overly fine-grained interleaving of stage partitioning and input sequence partitioning, which also implies more communication calls in Tensor Parallelism (although the total communication volume remains unchanged), potentially leading to a decrease in performance. Another observation is that the efficiency of Seq1F1B becomes more pronounced as the sequence length increases. This is because the computation time for each micro-sequence extends with longer sequences, thereby enhancing the benefits derived from sequence partitioning.
@@ 5 Conclusion
Table 6: The Ablation experiments based on 2.7B GPT of sequence partitioning strategies, where “w/o cwp” indicates the absence of a computation-wise partitioning strategy.
@@
Seq1F1B-I 122.7±0.0 1.18x
In this paper, we present Seq1F1B, an efficient 1F1B pipeline parallel scheduling method orienting to training Transformer-based LLMs on long sequences by decomposing the batch-level schedulable units used by typical 1F1B methods into more fine-grained sequence-level units. To achieve a better workload balance of the sequence-level pipeline, we design a computation-wise sequence partition strategy to partition the sequences well. Meanwhile Seq1F1B can integrate with other pipeline parallel methods such as 1F1B with interleaved stage or zero-bubble-pipeline. Our evaluations demonstrate that Seq1F1B outperforms the 1F1B and 1F1B-I scheduling strategies regarding memory efficiency and training throughput under variable sequence lengths and model sizes. Moreover, Seq1F1B can support the efficient training of a 30B GPT model on sequences up to 64k in length using 64xA100 GPUs, without using recomputation strategies, which is unachievable with existing pipeline parallel methods. In the future, we will thoroughly combine our method with other distributed methods to achieve better LLM training acceleration. In addition, we will systematically release our code to support the community in training LLMs to process longer sequences more efficiently.

### v5

@@ 3.2 Framework of Seq1F1B
where $P$ is the size of the PP and $k$ indicates the number of divisions of the sequence. This equation ensures that the last stage can perform a backward pass on the last sub-sequence of the first micro-batch when entering the steady phase, and the device responsible for each stage performs one more forward pass than the device responsible for the subsequent stage. Here, we construct a partially ordered queue $Q_{s}$, where each pop returns the tail sequence from the earliest enqueued intermediate states. This satisfies the FIFO principle in the batch dimension and the first-in-last-out (FILO) principle in the sequence dimension. In each step of the warm-up phase, devices execute one forward pass and enqueue the corresponding intermediate states of sub-sequences into $Q_{s}$. During the steady phase, after each device completes a forward pass, it dequeues intermediate states from $Q_{s}$ and performs a backward pass, following the standard 1F1B process, except that the units for forward and backward passes become a sub-sequence. During the cooling-down phase, devices dequeue the remaining intermediate states from $Q_{s}$, perform backward passes for remaining subsequences and accumulate the gradient to ensure mathematical equivalent with other synchronous schedules.
@@
From the timeline shown in Figure 1, it is evident that the Seq1F1B schedule offers a shorter execution time and significantly fewer bubbles, compared to the original 1F1B schedule. Meanwhile, it can be seen that each device now has less memory consumption since the sub-sequence is smaller than the micro-batch. Another observation is that optimizations similar to ZB1P can also be applied to Seq1F1B by delaying the gradient computation associated with weights in the backward pass. For more details, we refer to our appendix.
@@
Table 1: Settings used in experiments for training LLMs.
3.3 Framework of Seq1F1B-I
As shown in Figure 2, 1F1B-I (Narayanan et al., 2021b) achieves better efficiency by modifying the 1F1B schedule to support interleaved stages among devices. In 1F1B-I, each device is assigned multiple stages. Suppose we have $P$ devices and $V$ stages $\{s_{1},s_{2},\ldots,s_{V}\}$ in our pipeline, where $V$ is a multiple of $P$. The $i$-th device will handle $n$ stages $\{s_{i},s_{i+P},s_{i+2P},\ldots,s_{i+(n-1)P}\}$, where $n=\frac{V}{P}$. The number of warm-up micro-batches of each device $i$ in 1F1B-I is as follows,
$w_{i}=(P-i)\times 2+(n-1)\times P,i\in[1,P],$ (5)
After completing $P$ iterations of forward and backward passes, each device switches its context to the next stage that the device is responsible for. From Figure 2, the above part shows a 1F1B-I pipeline with $P=4$ and $V=8$, in which each device handles 2 stages. The 1F1B-I schedule reduces the bubble ratio by interleaving stages among devices. However, this interleaving slightly increases memory consumption, as the number of warm-up micro-batches $w_{i}$ is greater than that of 1F1B.
Similar to 1F1B-I, Seq1F1B-I further modifies 1F1B-I to achieve a sequence-level schedule. From Figure 2, Seq1F1B-I effectively reduces pipeline bubbles and the memory footprint of intermediate states compared to 1F1B-I. Seq1F1B-I defines the number of warm-up sub-sequences as
$w_{i}=(P-i)\times 2+(n-1)\times P+k-1,i\in[1,P],$ (6)
where $P$ is the size of the PP and $k$ indicates the number of divisions of the sequence. Using the partially ordered queue, Seq1F1B-I maintains a strict order of forward and backward passes as well as ensures the consistent semantics of gradient updates. From the perspective of pipeline bubbles, Seq1F1B-I outperforms both Seq1F1B and 1F1B-I. Besides, Seq1F1B-I requires slightly more memory than Seq1F1B but significantly less than 1F1B-I.
3.4 Workload Balance
In this section, we detail the strategy of sequence partition and workload balance consideration. Previous works, such as  (Li et al., 2021), have discussed strategies for sequence partitioning. To achieve an efficient pipeline schedule, the processing cost for each sub-sequence must be approximately equal to avoid pipeline bubbles. To this end, we design a computation-wise partition strategy by estimating the FLOPs of sequences and constructing a theoretical solution aiming to make the FLOPs of all sub-sequences as closely as possible. For a input sequence $S=\{x_{1},x_{2},\cdots,x_{n}\}$, we devide it into $k$ segments $S=\{S_{1},S_{2},\cdots,S_{k}\}$. Each segment has a length of $n_{i}$, where $\sum_{i=1}^{k}n_{i}=n$. We expect the computational amount of each segment to be roughly the same, that is
$\text{FLOPs}(S_{1})$ $=\text{FLOPs}(S_{2})$ (7)
$=\cdots=\text{FLOPs}(S_{k})$
$={\text{FLOPs}(S)\over k}.\vspace{-2em}$
Specifically, we use the method proposed in (Hoffmann et al., 2022) to estimate the FLOPs for each subsequence, formulated as
$\text{FLOPs}(S_{i})$ $=2n_{i}P+2Ln_{i}\left(\sum_{j=0}^{i}n_{j}\right)d,\forall i\in[1,k],$ (8)
$\text{FLOPs}(S)$ $=2nP+2Ln^{2}d,$
in which, $L$ is the number of layers, $d$ is the dimension of the model, and $P$ is the total number of parameters in the model. We have $k$ variables in Eq. (8) and $k$ equations in Eq. (7), and thus we can set up the equation to get the optimal segmentation.
Figure 3: Peak Memory consumption of training a series of models under varying sequence lengths and fixed batch settings. “X” means experiments ran out of memory. We take the maximum memory consumption between all devices for better clarification.
@@ 3.4 Workload Balance
Table 2: 2.7B GPT training experiments with PP size of 8 under 8$\times$ A100 setting.
@@
Table 3: 7B GPT training experiments with PP size of 4 and TP size of 8 under 32$\times$×A100 setting.
4 Experiments
4.1 Experimental Settings
In experiments, we measure Seq1F1B, Seq1F1B-I, 1F1B, and 1F1B-I under variable sequence lengths, different numbers of micro-batches, different numbers of GPUs, and different PP and TP sizes. Compared methods are as follows:
(1) Seq1F1B: Seq1F1B with computation-wise sequence partition strategy.
(2) Seq1F1B-I: Seq1F1B with interleaved stages and computation-wise sequence partition strategy.
(3) 1F1B/1F1B-I: 1F1B and 1F1B with interleaved stages in Megatron implementation.
(4) Seq1F1B w/o cwp: Seq1F1B without computation-wise sequence partition strategy.
(5) Seq1F1B-I w/o cwp: Seq1F1B-I without computation-wise sequence partition strategy.
All assessments are based on the GPT model and model configurations are listed in Table 1. All experiments focus on long-sequence training since a lot of work has mentioned its importance. For hyperparameter configurations, we set the number of sequence splits to four and each device manages two stages in interleaved settings. Our implementation is based on the open-source Megatron-LM project (Narayanan et al., 2021b) and ensures reproducibility. We adopt Megatron-V3 (Korthikanti et al., 2023)’s tensor parallelism in all experiments since it is necessary for long sequence training.
4.2 Main Results
@@
Table 4: 13B GPT training experiments with PP size of 4 and TP size of 8 under 32$\times$ A100 setting.
@@
Table 5: 30B GPT training experiments with PP size of 8 and TP size of 8 under 64$\times$ A100 setting.
In Figure 3, we compared the memory consumption of our method with that of 1F1B and 1F1B-I. As can be seen, our method consistently requires less memory across all settings. Notably, it can support training a 30B model on a 64$\times$ A100 cluster, which is impossible for the traditional combination of PP and TP. Additionally, we recorded TFLOPS (teraFLOPS) per GPU in our experiments to measure the hardware utilization of different methods. From Table 2, 3, 4 and 5, our method Seq1F1B outperforms 1F1B and 1F1B-I under almost all settings in both training throughput and teraFLOPS.
However, as observed in Table 3, 4, and 5, the Seq1F1B-I may have a performance degradation under multi-node settings. This could be due to the overly fine-grained interleaving of stage partitioning and input sequence partitioning, which also implies that more communication calls in TP (although the total communication volume remains unchanged), potentially leads to a decrease in performance. Another observation is that the efficiency of Seq1F1B becomes more pronounced as the sequence length increases. This is because the computation time for each micro-sequence extends with longer sequences, thereby enhancing the benefits derived from sequence partitioning.
@@ 5 Conclusion
Seq1F1B-I 122.7±0.0 1.18$\times$
Table 6: The Ablation experiments are based on 2.7B GPT of sequence partitioning strategies, where “w/o cwp” indicates the absence of a computation-wise partitioning strategy.
In this paper, we present Seq1F1B, an efficient 1F1B pipeline parallel schedule orienting to training Transformer-based LLMs on long sequences by decomposing the batch-level schedulable units used by typical 1F1B methods into more fine-grained sequence-level units. To achieve a better workload balance of the sequence-level pipeline, we design a computation-wise sequence partition strategy to partition the sequences well. Meanwhile, Seq1F1B can integrate with other pipeline parallel methods such as 1F1B with interleaved stage or zero-bubble-pipeline. Our evaluations demonstrate that Seq1F1B outperforms the 1F1B and 1F1B-I schedules regarding memory efficiency and training throughput under variable sequence lengths and model sizes. Moreover, Seq1F1B can support the efficient training of a 30B GPT model on sequences up to 64k in length using 64$\times$ A100 GPUs, without recomputation strategies, which is unachievable with existing pipeline parallel methods. In the future, we will thoroughly combine our method with other distributed methods to achieve better LLM training acceleration. In addition, we will systematically release our code to support the community in training LLMs to process longer sequences more efficiently.
Limitations
The current implementation of Seq1F1B is optimized for long-context training in LLMs, which may result in performance degradation when dealing with short context such as 4k/8k. We recommend using Seq1F1B in environments with limited communication bandwidth, as the PP incurs fewer communication costs compared to other parallel strategies.
Appendix A Appendix
A.1 Integration with Zero-bubble-pipeline
From Figure 4, we can see that Seq1F1B can integrate with the ZB1P method and further reduce bubbles while reducing memory demands by splitting sequence. Such integration outperforms simple ZB1P in both memory demands and pipeline bubbles since sequence-level pipelines naturally have fewer bubbles. Furthermore, Seq1F1B can integrate with ZB2P and ZBV methods too. Theoretically, introducing a zero-bubble-pipeline to Seq1F1B should be more efficient. Even though, such a fine-grained handcraft schedule may have performance degradation under some settings. We hope our work inspires future work to solve this problem.